### v1

Abstract
We review the literature on boundary discontinuity (BD) designs, a powerful non-experimental research methodology that identifies causal effects by exploiting a thresholding treatment assignment rule based on a bivariate score and a boundary curve. This methodology generalizes standard regression discontinuity designs based on a univariate score and scalar cutoff, and has specific challenges and features related to its multi-dimensional nature. We synthesize the empirical literature by systematically reviewing over $80$ empirical papers, tracing the method’s application from its formative uses to its implementation in modern research. In addition to the empirical survey, we overview the latest methodological results on identification, estimation and inference for the analysis of BD designs, and offer recommendations for practice.
Keywords: regression discontinuity, treatment effects estimation, causal inference.
Contents
1 Introduction
The Regression Discontinuity (RD) design is one of the leading observational methods for program evaluation and causal inference (see Abadie-Cattaneo_2018_ARE; Hernan-Robins_2020_Book, for overviews and further references). In its canonical form, a binary treatment is assigned to units whose value of a score is equal to or above a known cutoff, and not assigned to units whose score is below the cutoff. Under the key assumption that all observable and unobservable pretreatment characteristics vary smoothly around the cutoff, the discontinuous change generated by the hard-thresholding treatment assignment rule based on the univariate score and scalar cutoff can be used to learn about causal treatment effects. Lee-Lemieux_2010_JEL offer an early review with a list of standard RD empirical applications, Cattaneo-Titiunik_2022_ARE give a recent overview of the RD methodological literature, and Cattaneo-Idrobo-Titiunik_2020_CUP; Cattaneo-Idrobo-Titiunik_2024_CUP provide a practical introduction to modern RD methods and practice.
The Boundary Discontinuity (BD) design generalizes the RD design to the case where the score is bivariate, and the treatment is assigned to units according to the location of their score relative to a known boundary that splits its support into two disjoint regions. This setup is sometimes called the Multi-Score RD design (Papay-Willett-Murnane_2011_JoE; Reardon-Robinson_2012_JREE; Wong-Steiner-Cook_2013_JEBS), or the Geographic RD design (Keele-Titiunik_2015_PA; Keele-Titiunik-Zubizarreta_2015_JRSSA; Keele-Titiunik_2016_PSRM; Keele-etal_2017_AIE; Galiani-McEwan-Quistorff_2017_AIE; Rischard-Branson-Miratrix-Bornn_2021_JASA; Diaz-Zubizarreta_2023_AOAS). In this chapter, we discuss the empirical strategies most commonly employed for the analysis and interpretation of BD designs, and review the most recent theoretical and methodological results characterizing their econometric properties (Cattaneo-Titiunik-Yu_2025_BDD-Distance; Cattaneo-Titiunik-Yu_2025_BDD-Location; Cattaneo-Titiunik-Yu_2025_BDD-Pooling).
We introduce the BD design formally in Section 2, and then outline our review of the empirical literature in Economics, Political Science, Education, and other disciplines. Table 5 reports over $80$ empirical papers, which form the basis for the organization of this chapter. We find that the overwhelming majority of papers collapse the bivariate score into a univariate distance score measuring the shortest distance to the treatment assignment boundary, and then report a single average treatment effect estimated by pooling all the observations that are close to the boundary, regardless of their specific location. In contrast, only a handful of empirical papers investigate heterogeneity by localizing to specific regions or points along the boundary, despite the potential importance of this rich information for policy evaluation and decision-making.
Following our review of the empirical literature, in Section 3 we discuss pooling-based methods. We begin by discussing the seminal work of Card-Krueger_1994_AER, who appear to be the first to employ the BD design conceptually, although they did not localize to the assignment boundary for estimation purposes. We then discuss holmes1998effect and black1999better, who appear to be the first two papers to employ localization to the assignment boundary explicitly as their main empirical strategy. Subsequent empirical work has recognized the importance of localization to the boundary, but has considered different empirical strategies for treatment effect estimation via local flexible regression methods. One of the most influential papers is dell2010persistent, who incorporated a polynomial expansion of the bivariate location score as part of the local regression specification. More recently, an alternative empirical approach views the estimation based on distance to the boundary as a pooled univariate RD design, and thus estimates treatment effects using a local regression including a polynomial expansion of the univariate distance score. Recent examples include ito2020willingness and dehdari2022origins. Many empirical papers also include boundary-segment fixed effects in their local regression specifications, and some include interactions with the treatment indicator and/or between the polynomial expansions of the univariate distance score and the bivariate location score. Furthermore, for simplicity, or because distinct treatments occur in different regions of the boundary, some papers reduce the BD design to boundary-segment-specific univariate RD designs—see ou2010leave, LondonoVelezRodriguezSanchez_2020_AEJ and salti2022impact. Putting all together, Section 3 discusses eight distinct pooled local regression specifications commonly found in empirical work.
Building on our review of empirical strategies used for pooled estimation in BD designs, in Section 3.1 we also overview the main methodological results in Cattaneo-Titiunik-Yu_2025_BDD-Pooling for pooling-based estimation and inference. These methods are viewed as univariate RD designs based on the shortest distance to the assignment boundary, and hence are related to identification, estimation, and inference over a specific one-dimensional submanifold curve on the plane, defined by the assignment boundary. Recognizing this mathematical structure, Cattaneo-Titiunik-Yu_2025_BDD-Pooling establish their formal results leveraging concepts from geometric measure theory (federer2014geometric). After reviewing these ideas, we briefly discuss how estimation and inference can be conducted using the state-of-the-art methods from the RD design literature (Calonico-Cattaneo-Titiunik_2014_ECMA; Calonico-Cattaneo-Farrell-Titiunik_2019_RESTAT; Calonico-Cattaneo-Farrell_2020_ECTJ). Some of this work is conceptually related to a recent paper by chen2025semiparametric, who study estimation and inference for integral functionals on submanifolds when using nonparametric sieve estimation.
Section 4 discusses heterogeneity analysis and subsequent aggregation of causal treatment effects along the boundary. Although the boundary average treatment effect parameter emerging from pooling-based methods discussed in Section 3 is often of interest, and provides a useful aggregate measure of average treatment effects, the richness of the two-dimensional assignment in BD designs offer the opportunity to explore heterogeneous treatment effects along the assignment boundary. In geographic applications, this heterogeneity is directly linked to the geographic location of the units, providing useful information about how treatment effects vary in space; in non-geographic settings, the heterogeneity is directly linked to features of the units that are captured by the bivariate score variable, and hence the interpretation is necessarily application-specific. Thus, Section 4 begins by introducing the boundary average treatment effect curve, which captures the average treatment effect for each point along the assignment boundary. We also discuss how other aggregate causal effects can be recovered from that building block parameter, such as the weighted boundary average treatment effect and the largest average treatment effect.
After introducing the main ideas and causal parameter of interest, Section 4 reviews the theoretical and methodological takeaways from Cattaneo-Titiunik-Yu_2025_BDD-Distance and Cattaneo-Titiunik-Yu_2025_BDD-Location. These two recent papers investigate identification, estimation, and inference for heterogeneous treatment effects along the assignment boundary, and transformations thereof, from two connected but distinct perspectives: Cattaneo-Titiunik-Yu_2025_BDD-Distance studies boundary-point specific distance-based methods, which are directly motivated by the poloing-based approaches discussed in Section 3, while Cattaneo-Titiunik-Yu_2025_BDD-Location studies methods based on the bivariate location directly. For each approach, the estimation and inference results are valid both pointwise and uniformly along the assignment boundary, in addition to providing the basis for constructing other treatment effect estimators.
Finally, Section 5 offers recommendations for practice, and concludes. The methods discussed in this chapter can be implemented using the general-purpose software packages rdrobust, rdhte, rdmulti, and rd2d, located at https://rdpackages.github.io/, where replication files and related references are also available.
2 The Boundary Discontinuity Design
Each unit $i$ in the study, $i=1,\ldots,n$, has a continuous bivariate score $\mathbf{X}_{i}=(X_{1i},X_{2i})$ that takes values in the set $\mathcal{X}\subseteq\mathbb{R}^{2}$. The assignment of units to the treatment or control condition depends on the location of their score $\mathbf{X}_{i}$ relative to a known one-dimensional boundary curve $\mathcal{B}$ that splits $\mathcal{X}$ into two disjoint regions: $\mathcal{X}=\mathcal{A}_{0}\cup\mathcal{A}_{1}$, with $\mathcal{A}_{0}$ and $\mathcal{A}_{1}$ the control and treatment disjoint (connected) regions, respectively, and $\mathcal{B}=\mathtt{bd}(\mathcal{A}_{0})\cap\mathtt{bd}(\mathcal{A}_{1})$, where $\mathtt{bd}(\mathcal{A}_{t})$ denotes the boundary of the set $\mathcal{A}_{t}$. (We follow convention and assume that the boundary belongs to the treatment group, that is, $\mathtt{bd}(\mathcal{A}_{1})\subset\mathcal{A}_{1}$ and $\mathcal{B}\cap\mathcal{A}_{0}=\emptyset$.) The observed outcome variable is $Y_{i}$.
An ancestor of the BD design was used by Card-Krueger_1994_AER in their seminal study of the effects of the minimum wage on employment. In their analysis, the treatment of interest is the increase of the state minimum wage in New Jersey that went into effect on April 1, 1992, and rose the minimum wage to \$5.05. The authors compared a sample of fast food restaurants in New Jersey to a sample of fast food restaurants in eastern Pennsylvania, a state that shares a border with New Jersey and where the minimum wage did not increase. After the increase in New Jersey, the average starting wage was \$5.08 in the New Jersey sample and \$4.62 in the Pennsylvania sample.
The rationale for comparing restaurants in New Jersey to restaurants in eastern Pennsylvania was that, by focusing on close geographic areas, these restaurants would be subject to similar economic conditions that would otherwise confound the effects of changes to the minimum wage. Although Card-Krueger_1994_AER also relied on a difference-in-differences design comparing both samples before and after the increase, they explicitly mentioned the close geographic proximity between the areas to justify their choice of comparison group, as they believed that fast-food stores in eastern Pennsylvania formed a “natural basis for comparison with the experiences of restaurants in New Jersey” due to Pennsylvania being a “nearby state[s]” (Card-Krueger_1994_AER, p. 773).
The use of close proximity as the basis for reducing potential confounders and enhancing the credibility of causal interpretations bears a direct connection to the BD design. The main difference is that Card-Krueger_1994_AER included non-adjacent areas in New Jersey and Pennsylvania (e.g., the New Jersey sample included restaurants in the New Jersey shore, far away from the Pennsylvania border), and did not directly use the unit’s distance to the boundary in their analysis. The explicit use of distance to the assignment boundary is central to the BD design, which is why we consider Card-Krueger_1994_AER to be an ancestor rather than an instance of the BD design itself: while the authors leverage the change in treatment assignment induced by a geographic boundary, they did not explicitly localize to the boundary itself.
holmes1998effect and black1999better are the earliest empirical examples that we could find employing the BD design. holmes1998effect used a BD design to study the effect of pro-business policies on manufacturing, comparing states that adopted right-to-work laws with adjacent states that did not. Using the longitude and latitude coordinates of each county’s centroid, he calculated the minimum distance of the centroid to the assignment border, and focused the analysis on counties close to the border as a way to “control for differences across states in these various characteristics that are unrelated to policy” (holmes1998effect, p. 671). Similarly, black1999better used a BD design to infer the quality of public schools from housing prices. She compared house prices on opposite sides of school attendance district boundaries, restricting the sample to houses close to the boundary. The justification for this localization was that houses on opposite sides of the boundary within a small area around it would be similar in all characteristics except for school quality—as the latter changes discontinuously at the attendance district boundary. In these examples, the bivariate score is the pair of latitude and longitude coordinates that determines the location of each unit (county or house), and the assignment boundary is the geographic border between two adjacent states or school attendance districts; the treatment is the policy or feature that changes abruptly at the border—labor policy for holmes1998effect and school quality for black1999better. This type of BD design is often referred to as a Geographic RD design.
One of the most common non-geographic applications of the BD design occurs in education when a treatment is given on the basis of two exam scores. For example, ou2010leave studies the effect of a high school exit exam on the likelihood of dropping out of high school. The exam has a mathematics and a language arts component, each of which is graded separately; students must achieve a minimum proficiency in each section in order to pass the exam. The bivariate score in this case is the pair of mathematics and language test scores, and the treatment is passing the exam.
Figure 1 illustrates the boundary and treatment assignment regions in two stylized settings corresponding to the geographic and non-geographic examples. Figure 1(a) illustrates the geographic BD design case, where units are split into adjacent treated and control areas according to their location with respect to a geographic border, as in black1999better and holmes1998effect. (This figure shows the New Jersey-Pennsylvania border as considered by Card-Krueger_1994_AER.) Figure 1(b) illustrates the non-geographic setting, using as example the study by ou2010leave where students receive two exam scores and a treatment is given only to those individuals who score above a minimum cutoff in each exam. In geographic applications, the bivariate score is always composed of the geographic coordinates of the units. In non-geographic applications, the bivariate score could include different types of variables; in addition to test scores from multiple exams, non-geographic bivariate scores include shares of airline passengers in origin and destination cities (snider2015barriers), tax rates and income ratios (egger2015impact), systolic and diastolic blood pressure measurements (dai2022effects), different component scores in a means testing formula (salti2022impact; kampfen2024heterogeneous), and population counts at different time periods (hinnerich2014democracy).
(a) Geographic BD Design
(b) Non-geographic BD Design
Figure 1: BD Design Illustration
Figure 1 illustrates an important distinction between the geographic and non-geographic examples. In non-geographic examples, the function that describes the boundary tends to be known, and is typically linear with few irregularities and very few kink points. In contrast, in geographic BD designs the boundary is typically the border between political or administrative subdivisions; as a consequence, its shape is only available as coordinates on a map, and its geometry can be complicated. In the best case possible, it will be piecewise linear with many kink points, but it will be more irregular in many applications. In his early study, holmes1998effect already recognized this challenge and discussed explicitly the complex boundary shapes created by geographic borders. Indeed, since the influential work by mandelbrot1967long; mandelbrot1983fractal, there has been an ongoing debate among mathematicians and philosophers trying to decide whether geographic borders (and other shapes in nature) are fractals (see avnir1998geometry, and references therein). From a methodological perspective, as recently formalized in Cattaneo-Titiunik-Yu_2025_BDD-Distance; Cattaneo-Titiunik-Yu_2025_BDD-Location; Cattaneo-Titiunik-Yu_2025_BDD-Pooling, the geometry of the assignment boundary in BD designs has fundamental implications for identification, estimation, and inference. We will review some of these recent findings in the upcoming sections.
In the standard RD design, the distance between the univariate score and the single cutoff is naturally always measured by the Euclidean distance, and observations are “close” to the cutoff $c$ when their univariate score $X_{i}$ implies a “small” distance $|X_{i}-c|$. In contrast, in the BD design, “closeness” to the boundary can be defined in different ways, depending on whether the parameter of interest is defined relative to a specific point on the boundary, or as some aggregate thereof along the entire boundary, such as an average or supremum. Furthermore, due its multivariate nature, the analysis can depend on the specific notion of distance used.
In the next sections, we discuss the two main approaches to the analysis of BD designs. The first approach uses the variation in treatment assignment induced by entire boundary simultaneously, and thus pools all observations sufficiently “close” the boundary in a single analysis, producing a single average treatment effect across all boundary points. This pooling approach relies on localization to $\mathcal{B}$ by first computing the distance between each observation’s location and the nearest point on the boundary, denoted by $D_{i}$, and then retaining only those observations for which this distance is no larger than a specific bandwidth, denoted by $h$. The advantage of this approach is that it naturally mimics the univariate RD design, using the bivariate location of each unit to build a univariate score that is used directly in a “standard” univariate RD analysis. This approach thus produces a single average treatment effect estimate.
In contrast, a second approach focuses on estimating the average treatment effect at each point on the boundary, performing localization relative to the specific point of analysis, thereby retaining observations whose distance to that point is no greater than a chosen bandwidth. Estimation and inference are conducted separately at each boundary point $\mathbf{x}\in\mathcal{B}$, which captures the heterogeneity of treatment effects along the boundary. Aggregate causal effects can then be recovered by averaging (or applying other transformations to) these boundary-point specific average treatment effects, including as a special case the average treatment effect along the entire boundary that is the focus of the pooling approach. The boundary-point-specific approach can be seen as focusing on heterogeneous average treatment effects that can then be used as building blocks to construct more aggregated causal parameters. The advantage of this approach is that it captures the full richness of the BD design by first learning the heterogeneity of the average treatment effect along the boundary, which usually summarizes unit characteristics of high policy relevance such as geographic location, academic performance, or economic need.
2.1 BD Designs in Practice
We searched the academic literature in multiple social sciences to learn how applied researchers implemented the BD design. We conducted our search in four general queries. In our initial Economics query, we included ten leading journals in Economics—Journal of Political Economy, Review of Economic Studies, Econometrica, American Economic Review, Quarterly Journal of Economics, Review of Economics and Statistics, American Economic Journal: Applied Economics, American Economic Journal: Economic Policy, American Economic Journal: Macroeconomics, and American Economic Journal: Microeconomics. In our initial Political Science query, we searched six leading Political Science journals—American Political Science Review, American Journal of Political Science, Journal of Politics, Political Science Research and Methods, British Journal of Political Science, and Quarterly Journal of Political Science. In both cases we searched for terms like “boundary discontinuity”, “multivariate regression discontinuity”, and “multiscore regression discontinuity”, and focused our attention on papers published since $2015$. We also expanded our search to include other articles that were repeatedly cited by the articles collected in our initial queries.
We also searched in the field of Education, as methodological researchers in that area were among the earliest scholars who discussed the multi-score RD design in non-geographic settings. For this search, we started with the seminal methodological papers in Education by Reardon-Robinson_2012_JREE and Wong-Steiner-Cook_2013_JEBS, and collected papers that both cited this methodological papers and were cited by them. Finally, we also included some papers in environmental science, to illustrate uses of the design at the intersection between the natural and social sciences. Our review is not meant to provide an exhaustive list of the literature. This literature is vast, and cannot be fully captured by our approach since we excluded several journals and restricted the time period. Our goal is simply to provide an overview of recent empirical work using the BD design across the social sciences to gauge the most common strategies used for empirical analysis.
The results of our search are summarized in Table 5. Our search yielded 82 published papers. Of these, we found that the overwhelming majority (76 articles or approximately 93%) used a pooling approach, focusing on the average effect along the entire boundary and using the univariate closest distance to the boundary as the running variable. Fewer than 10 articles were explicit in accounting for the heterogeneity along the boundary; of these, only 5 reported effects for different points or segments along the boundary (gonzalez2021cell; grout2011land; Keele-Titiunik_2015_PA; snider2015barriers; velez2019tuning). In sum, our review of the practical literature indicates that most applied researchers adopt a pooling approach, performing localization by grouping all observations sufficiently “close” to the boundary and then estimating a single average treatment effect. These common practices motivate the taxonomy we used in the rest of this chapter.
3 Pooling-Based Methods
Given a sample of outcomes and bivariate scores, $(Y_{1},\mathbf{X}_{1}),\ldots,(Y_{n},\mathbf{X}_{n})$, a distance function $\mathcal{d}(\cdot,\cdot)$ is used to measure the closeness between any two points in $\mathcal{X}$. The most common example is the Euclidean distance $\mathcal{d}(\mathbf{x}_{1},\mathbf{x}_{2})=\|\mathbf{x}_{1}-\mathbf{x}_{2}\|=\sqrt{(x_{11}-x_{21})^{2}+(x_{12}-x_{22})^{2}}$ for $\mathbf{x}_{j}=(x_{j1},x_{j2})$, $j=1,2$, although in some applications this is not the most appropriate measure. For example, ambrus2020loss are interested in how close residences in a city are to each other in terms of how long it takes for a person to walk between them; this requires calculating a walking distance, which is different from the Euclidean distance which would not take into account, say, that the path must follow city blocks and cannot go through buildings. More broadly, in geographic settings it is common practice to employ specific geodetic distances (Banerjee_2005_Biometrics).
Once a distance function $\mathcal{d}(\mathbf{X}_{i},\mathbf{x})$ has been chosen, researchers who employ the pooling approach define the closest signed distance to the boundary,
$D_{i}\equiv\big(\mathds{1}(\mathbf{X}_{i}\in\mathcal{A}_{1})-\mathds{1}(\mathbf{X}_{i}\in\mathcal{A}_{0})\big)\cdot\inf_{\mathbf{x}\in\mathcal{B}}\mathcal{d}(\mathbf{X}_{i},\mathbf{x}),$
for $i=1,\ldots,n$. For a unit $i$, $D_{i}$ measures how far $i$’s location $\mathbf{X}_{i}$ is to the boundary point that is closest to $\mathbf{X}_{i}$, regardless of where that boundary point is located. Following the standard RD design logic, but now expanded to the BD design (a bivariate setting), localization occurs simultaneously along the entire assignment boundary because the univariate distance $D_{i}$ to $\mathcal{B}$ is used to construct the region such that $|D_{i}|\leq h$, which covers the boundary $\mathcal{B}$, where $h$ denotes a bandwidth parameter. This region, called a “tubular neighborhood” in mathematics (federer2014geometric), is a natural generalization of the standard localization approach used in univariate RD designs, where the region is simply $|X_{i}-c|\leq h$ when the score $X_{i}$ is scalar and $c$ is the univariate cutoff.
Figure 2 illustrates the idea graphically. In both cases, the tubular neighborhood represents the area that includes the observations used in the pooling-based analysis. Figure 2(a) shows a “large” bandwidth $h_{1}$, while Figure 2(b) shows a “small” bandwidth $h_{2}$. All units contribute simultaneously, regardless of their specific location. In particular, two units $i$ and $j$ with the same signed distance to the boundary, $D_{i}=D_{j}=d$, will be assigned to the same group (treatment or control, depending on the sign of $d$) and will be $|d|$ units away from the boundary, but could be arbitrarily far from each other.
(a) Large Bandwidth
(b) Small Bandwidth
Figure 2: Pooling-Based Method: Localization around Entire Assignment Boundary
Figure 3 presents the two resulting RD plots (Calonico-Cattaneo-Titiunik_2015_JASA) in the pooling approach, based on the outcome data $Y_{i}$ and the univariate distance score $D_{i}$ (the cutoff is $c=0$ by construction of $D_{i}$). These plots include a fourth-order global polynomial fit (solid line) for visualization purposes only; those global approximations are not recommended for estimation of treatment effects because they tend to exhibit poor performance due to the Runge’s phenomenon. Instead, as anticipated in Figure 2 and demonstrated in Figure 3, the idea is to localize around the “cutoff” (i.e., the one-dimensional boundary curve) to conduct estimation and inference.
(a) Large Bandwidth
(b) Small Bandwidth
Figure 3: Pooling-Based Method: RD Plots (with Global Quartic Polynomial Estimates)
In the pooling approach, the analysis is conducted using all the observations with score $\mathbf{X}_{i}$ within the tubular neighborhood determined by the bandwidth $h$, that is, for all units with $|D_{i}|\leq h$. However, for this subsample, the specific treatment effect estimation approach can vary substantially. To describe and unify the most common approaches in the empirical literature illustrated in Table 5 as well as other approaches recommended from a methodological perspective, we define the following generic weighted least squares regression notation:
$\mathtt{reg}\quad Y_{i}\quad\mathtt{on}\quad\mathbf{Z}_{i}^{\top}\qquad[\;W_{i}\;]\qquad\iff\qquad\min_{\boldsymbol{\beta}}\sum_{i=1}^{n}\big(Y_{i}-\mathbf{Z}_{i}^{\top}\boldsymbol{\beta}\big)^{2}\cdot W_{i}$
where $Y_{i}$ is the outcome variable, $\mathbf{Z}_{i}$ is a (column) vector of independent variables, and $W_{i}$ are weights. A localized regression analysis can be implemented by setting $W_{i}=\mathds{1}(|D_{i}|\leq h)$ or, more generally, $W_{i}=K(D_{i}/h)$ for a kernel function $K(\cdot)$ determining the relative weight given to each observation. The bandwidth $h$ determines the degree of localization, as illustrated in Figure 2. When $\mathbf{Z}_{i}$ contains polynomial expansions of $D_{i}$, the resulting estimation approach is often called nonparametric local polynomial regression but, as we will see shortly, empirical researchers often employ other variables in their local regression fit within the tubular determined by $D_{i}$.
We also define the treatment assignment indicator
$T_{i}=\mathds{1}(D_{i}\geq 0)=\mathds{1}(\mathbf{X}_{i}\in\mathcal{A}_{1}).$
All empirical approaches naturally include $T_{i}$ in $\mathbf{Z}_{i}$ to estimate treatment effects; the majority also include other variables such as $D_{i}$, $\mathbf{X}_{i}$, boundary-segment fixed effects (formally defined below), as well as certain transformations and interactions thereof. For simplicity, we omit discussing the inclusion of predetermined covariates, but we note that researchers often further augment the basic local regression specifications with pretreatment variables such as census characteristics, terrain ruggedness and elevation, time fixed effects, among other features. For a discussion of the role of preintervention covariates in the standard RD design see Calonico-Cattaneo-Farrell-Titiunik_2019_RESTAT for efficiency gains and Calonico-Cattaneo-Farrell-Palomba-Titiunik_2025_wp for treatment effect heterogeneity.
The pooling approach for analyzing BD designs begins by choosing the localization bandwidth $h$ to determine the units with $|D_{i}|\leq h$ that will be used in the subsequent estimation. Until now, there were no formal methods for choosing the bandwidth $h$ in an objective and data-driven way, so researchers often considered a range of “reasonable” values. For example, black1999better reported results in three different bandwidths, keeping observations within $0.35$, $0.25$, and $0.15$ miles of the nearest point on the boundary. As in standard univariate RD designs, different choices of $h$ trade off bias and variance in the estimation, with a larger $h$ leading to more bias but less variance, and a smaller $h$ having the opposite effect. Cattaneo-Titiunik-Yu_2025_BDD-Pooling recently established formal results characterizing this tradeoff and provided formal guidance for choosing the bandwidth when implementing pooling BD methods. We discuss these methods in Section 3.1.
Once the bandwidth $h$ has been chosen, researchers pool all units with distance $D_{i}$ within $h$ and then compare those assigned to the treatment group ($D_{i}\geq 0$) to those assigned to the control group ($D_{i}<0$). Because the bandwidth localizes the analysis to only a “small” region on either side of the boundary $\mathcal{B}$, the empirical analysis compares barely treated units to barely control units which, as in the standard univariate RD design, is the basis for the causal interpretation of the comparison, as all confounders are assumed to vary smoothly at the boundary while the treatment assignment changes abruptly from zero to one.
The most basic analysis proceeds by assuming that the units within the shrinking tubular of “width” $2h$ are as-if randomly assigned to treatment and control. This approach is akin to the “local randomization” framework in standard RD designs (Lee-Lemieux_2010_JEL; Cattaneo-Frandsen-Titiunik_2015_JCI; Cattaneo-Titiunik-VazquezBare_2017_JPAM). Thus, a natural treatment effect estimator is the difference-in-means of the outcome between treated and control units with distance to the boundary no larger than the bandwidth:
$\mathtt{reg}\quad Y_{i}\quad\mathtt{on}\quad 1,\;T_{i}\qquad[\;\mathds{1}(|D_{i}|\leq h)\;],$ (1)
where the coefficient on $T_{i}$ is
$\widehat{\tau}_{\mathtt{DIM}}=\frac{1}{N_{1}(h)}\sum_{i=1}^{n}\mathds{1}(0\leq D_{i}\leq h)\cdot Y_{i}-\frac{1}{N_{0}(h)}\sum_{i=1}^{n}\mathds{1}(-h\leq D_{i}<0)\cdot Y_{i},$
with $N_{0}(h)=\sum_{i=1}^{n}\mathds{1}(-h\leq D_{i}<0)$ and $N_{1}(h)=\sum_{i=1}^{n}\mathds{1}(0\leq D_{i}\leq h)$. This specification can be interpreted as a local-constant polynomial fit on $D_{i}$; it does not include the score in the fit because under the assumption of local randomization this score is uncorrelated with the outcome.
(a) Large Bandwidth
(b) Small Bandwidth
Figure 4: Pooling-Based Method: Local-Constant Estimation ($p=0$)
For example, black1999better reported results using specification (1), calling it a “rough, nonparametric estimate of the value of better schools” (p. 590). She also augmented the basic specification to include “segment fixed effects”, that is, binary variables indicating whether an observation is closest to one of several segments partitioning the assignment boundary. To describe this approach formally, we let $\mathcal{B}_{\ell}$, $\ell=1,\ldots,L$, be disjoint segments that partition the assignment boundary, that is, $\mathcal{B}=\sqcup_{1\leq\ell\leq L}\mathcal{B}_{\ell}$. Then, $S_{i}=\operatorname*{arg\,min}_{1\leq\ell\leq L}d(\mathbf{X}_{i},\mathcal{B}_{\ell})$ is the boundary-segment allocation random index for each unit $i=1,\ldots,n$, where it is understood that when the $\operatorname*{arg\,min}$ contains more than one element, the smallest item is chosen. The vector of segment fixed effects for observation $i$ is $\boldsymbol{\iota}_{L}(S_{i})=(\mathds{1}(S_{i}=1),\cdots,\mathds{1}(S_{i}=L))$, which collects the $L$ boundary segment binary indicators determined by $i$’s location. The difference-in-means specification with segment fixed effects is
$\mathtt{reg}\quad Y_{i}\quad\mathtt{on}\quad\boldsymbol{\iota}_{L}(S_{i}),\;T_{i}\qquad[\;\mathds{1}(|D_{i}|\leq h)\;],$ (2)
where the intercept is removed due to its perfect collinearity with the boundary-segment fixed effects. Setting $L=1$ reduces $\boldsymbol{\iota}_{L}(S_{i})$ to the common intercept in specification (1).
For example, black1999better used a modified version of specification (2) where she included the test scores of the school associated with observation $i$ instead of the binary treatment assignment indicator $T_{i}$. The causal interpretation of this specification rests on the assumption that, after localizing the analysis to include only observations near the boundary of school attendance districts, the only factor that can explain a difference in average house prices between the treated and control areas is school quality, because school quality changes discontinuously at the boundary while all other determinants of house prices are equal on average. In this sense, using test scores instead of the treatment assignment indicator in specifications (1) and (2) is in the spirit of an instrumental variables strategy where $T_{i}$ is used as an instrument for test scores (which proxy school quality), and the localization to a small bandwidth around the boundary ensures that units assigned to treatment and units assigned to control are similar in all other characteristics that might affect house prices.
The seminal research design used by black1999better is prior to the development of modern methods for RD analysis, which occurred predominantly during the $2000$s; see Cattaneo-Titiunik_2022_ARE for an overview. In more recent practice, the simple specifications (1) and (2) are not commonly used because researchers tend to adopt a continuity-based approach rather than a local randomization approach for interpretation. Under the continuity-based approach, the local constant fit has poor control of the misspecification bias, and hence practitioners prefer more flexible local regression specifications.
Later empirical work using BD designs made more explicit the connection between geographic discontinuities in treatment assignment and the classical RD design, following more closely the methods of univariate continuity-based RD analysis, which emphasized including the RD score in the polynomial specification to allow for arbitrary dependence between the RD score and outcome even within the bandwidth. Although there is considerable diversity in the specifications used in more recent practice, most researchers that include the score adopted one of two main strategies: a polynomial fit of the outcome on the bivariate score $\mathbf{X}_{i}$, or a polynomial fit of the outcome on the univariate distance $D_{i}$.
In an influential contribution, dell2010persistent augmented specification (2) by including the bivariate score $\mathbf{X}_{i}$. She studied the long-run impact of the mita, a forced mining labor system instituted by the Spanish in Peru and Bolivia between $1573$ and $1812$, using the historical boundary that determined which communities were forced to send labor to define treated and control areas. She studied the effect of the mita on contemporary outcomes measured at the household level, including consumption and childhood stunting. Her main local regression specification was
$\mathtt{reg}\quad Y_{i}\quad\mathtt{on}\quad\boldsymbol{\iota}_{L}(S_{i}),\;T_{i},\;\mathbf{r}_{p}(\mathbf{X}_{i})\qquad[\;\mathds{1}(|D_{i}|\leq h)\;],$ (3)
which includes the treatment indicator, boundary segment fixed effects, and a polynomial expansion of the bivariate score $\mathbf{X}_{i}$ (with $p=3$ in her case). The estimated coefficient of interest is always the coefficient that accompanies the treatment indicator $T_{i}$. In our notation, $\mathbf{r}_{p}(\mathbf{u})=(u_{1},u_{2},u_{1}^{2},u_{1}u_{2},u_{2}^{2},\ldots,u_{1}^{p},u_{1}^{p-1}u_{2},\ldots,u_{1}u_{2}^{p-1},u_{2}^{p})$ with $\mathbf{u}=(u_{1},u_{2})$. The inclusion of (a polynomial basis of) the bivariate location score $\mathbf{X}_{i}$ is meant to control for “the smooth effects of geographic location” dell2010persistent. A notable characteristic of this specification is that it does not include interactions between the bivariate score and the treatment indicator, which imposes the restriction that the polynomials be identical in the treated and control areas. As for bandwidth selection, dell2010persistent considered 100100km of the mita boundary first, but this choice is later reduced to 7575km and 5050km in subsequent analyses.
(a) Large Bandwidth
(b) Small Bandwidth
Figure 5: Pooling-Based Method: Local-Linear No-Interaction Estimation ($p=1$)
Specification (3) has become a common strategy to analyze BD designs. For example, mendez2022multinationals use it to study the effect of private sector companies in the development of local amenities, using a land concession to the multinational United Fruit Company in Costa Rica; paulsen2023foundations use it to estimate the effect of school funding on future earnings and political participation in New York, comparing areas that received higher funding to neighboring areas that did not; and de2018agents compare areas in South Africa where traditional leaders remain highly influential to areas where they are not, and estimate the electoral impact of strong political ties between traditional leaders and the African National Congress (ANC). Other examples are listed in Table 5. All of these examples follow specification (2) by including a flexible polynomial expansion of latitude and longitude, as well as boundary-segment fixed effects, but without allowing for interactions with the treatment indicator $T_{i}$.
An alternative approach is to directly mimic the standard univariate RD design, taking the scalar (signed) distance to the closest point on the boundary for every unit, $D_{i}$, as the univariate score, and employing local polynomial regression. The most basic local regression specification in this context takes the form:
$\mathtt{reg}\quad Y_{i}\quad\mathtt{on}\quad 1,\;T_{i},\;\mathbf{r}_{p}(D_{i})\qquad[\;\mathds{1}(|D_{i}|\leq h)\;],$ (4)
with the polynomial expansion reducing to the simpler vector of the univariate distance $\mathbf{r}_{p}(D_{i})=(D_{i},D_{i}^{2},\ldots,D_{i}^{p})$. As in the case of specification (3), specification (4) can also be augmented with boundary-segment fixed effects:
$\mathtt{reg}\quad Y_{i}\quad\mathtt{on}\quad\boldsymbol{\iota}_{L}(S_{i}),\;T_{i},\;\mathbf{r}_{p}(D_{i})\qquad[\;\mathds{1}(|D_{i}|\leq h)\;].$ (5)
In both cases, the treatment effect estimator is the coefficient estimate accompanying $T_{i}$. Figure 5 offers a graphical representation of this empirical strategy.
Although specifications (4) and (5) aim to mimic the univariate RD design, they differ from best practices in an important respect. In the RD design, the recommended specification is a polynomial fit of the outcome on an intercept, the treatment assignment indicator, a polynomial expansion of the score, and the interaction between this expansion and the treatment assignment indicator. The latter interaction is crucial to allow the polynomial fit for control units to be different from the polynomial fit for treated units. None of the specifications discussed so far allow for this flexibility, potentially introducing a larger misspecification bias.
The most flexible specification based on $D_{i}$ takes the form:
$\mathtt{reg}\quad Y_{i}\quad\mathtt{on}\quad\boldsymbol{\iota}_{L}(S_{i}),\;T_{i},\;\mathbf{r}_{p}(D_{i}),\;T_{i}\cdot\mathbf{r}_{p}(D_{i})\qquad[\;\mathds{1}(|D_{i}|\leq h)\;],$ (6)
where the estimated coefficient of interest continues to be the one accompanying $T_{i}$. In practice, specification (6) is often implemented with $p=1$ and, depending on whether standard software for RD design analysis is used, the specification may employ a non-uniform weighting kernel (e.g., a triangular kernel that down-weights observations as their distance from the boundary increases). Putting aside the inclusion of boundary-segment fixed effects, Figure 6 illustrates the importance of including the interaction term $T_{i}\cdot\mathbf{r}_{p}(D_{i})$ in the treatment effect estimation, as in specification (6). Comparing Figures 5 and 6 shows that the two approaches can lead to different local approximations, with specifications (1) through (5) potentially exhibiting a larger bias than specification (6).
(a) Large Bandwidth
(b) Small Bandwidth
Figure 6: Pooling-Based Method: Local-Linear Estimation ($p=1$)
Specification (6) is one of the most commonly used in recent practice. Examples include ayres2021environmental, eugster2017culture, jones2022factor, lalive2008extended, and lowes2021concessions. We note that several articles write their empirical specifications as (4) or (5), but they actually implement specification (6) in their analyses (as clarified in text, footnotes or replication files). These researchers employ standard practices and software from the univariate RD design literature, which follows specification (6).
Specification (3) can also be augmented to include the interaction term between the treatment assignment and the polynomial expansion of the bivariate score $\mathbf{X}_{i}$:
$\mathtt{reg}\quad Y_{i}\quad\mathtt{on}\quad\boldsymbol{\iota}_{L}(S_{i}),\;T_{i},\;\mathbf{r}_{p}(\mathbf{X}_{i}),\;T_{i}\cdot\mathbf{r}_{p}(\mathbf{X}_{i})\qquad[\;\mathds{1}(|D_{i}|\leq h)\;],$ (7)
where the estimated treatment effect coefficient accompanying $T_{i}$ enjoys better bias properties due to the added flexibility in the local regression specification. However, this more flexible approach is not as common in the empirical literature—some exceptions are gonzalez2021cell, egger2015impact, and jones2022effects.
In some applications, researchers estimate treatment effects in BD designs by considering two separate univariate RD designs, that is, keeping one dimension fixed and studying the other dimension of the bivariate score. For example, LondonoVelezRodriguezSanchez_2020_AEJ investigate the effects of Ser Pilo Paga (SPP), a government subsidy in Colombia that provided tuition support for post-secondary students to attend a government-certified higher education institution. Eligibility was based on both merit and economic need: in order to qualify for the program, students had to obtain a high grade in Colombia’s national standardized high school exit exam, SABER 11, and they also had to come from economically disadvantaged families, as measured by the survey-based wealth index SISBEN. The resulting BD design looks exactly as in Figure 1(b) with $\mathbf{X}_{i}=(X_{1i},X_{2i})=(\text{{SABER11}}_{i},\text{{SISBEN}}_{i})$. The authors analyzed it by considering each cutoff separately and then pooling all observations in the other dimension, leading to two univariate RD designs: one with score $\textit{SISBEN}_{i}$ in the sample $\text{{SABER11}}_{i}\geq 0$, and the other with score $\text{{SABER11}}_{i}$ in the sample $\textit{SISBEN}_{i}\geq 0$. This approach is equivalent to splitting the analysis along each of the two linear segments of $\mathcal{B}$ and then employing a pooling approach. This idea leads to the following local regression specification:
$\mathtt{reg}\quad Y_{i}\quad\mathtt{on}\quad T_{i}\cdot\boldsymbol{\iota}_{L}(S_{i}),\;\boldsymbol{\iota}_{L}(S_{i})\otimes\mathbf{r}_{p}(D_{i}),\;T_{i}\cdot\boldsymbol{\iota}_{L}(S_{i})\otimes\mathbf{r}_{p}(D_{i})\qquad[\;\mathds{1}(|D_{i}|\leq h)\;],$ (8)
where the indicator variable $T_{i}$ is removed due to the perfect collinearity with $T_{i}\cdot\boldsymbol{\iota}_{L}(S_{i})$, since $\boldsymbol{\iota}_{L}(S_{i})$ contains indicator variables capturing a partitioning of $\mathcal{B}$, and $\otimes$ denotes the Kronecker product. The treatment effect of interest is now the collection of coefficients accompanying $T_{i}\cdot\boldsymbol{\iota}_{L}(S_{i})$, one for each segment. Specification (8) is roughly equivalent to employing specification (6) $L$ times, in each case only using the data corresponding to one of $L$ segments. (The only minor difference is whether observations are assigned to only one segment, or reused for different segment-specific treatment effect estimation implementation.) For instance, ou2010leave, LondonoVelezRodriguezSanchez_2020_AEJ and salti2022impact employ this boundary-segment specific pooling-based approach with $L=2$, analyzing each of the two linear segments as in Figure 1(b). Another empirical example employing this approach is .
Finally, some researchers have also included both the univariate distance $D_{i}$ and the bivariate score $\mathbf{X}_{i}$ in their local regression specifications—see, for example, ehrlich2018persistent and dehdari2022origins. In addition, it is possible to conceptualize more flexible specifications by including other interaction terms between $T_{i}$, $\mathbf{r}_{p}(\mathbf{X}_{i})$, and $\boldsymbol{\iota}_{L}(S_{i})$ analogous to specification (8). These approaches are uncommon in practice according to our literature review (Table 5), so we do not discuss them further.
3.1 Methodological Results
While there has been a proliferation of different empirical strategies relaying on the pooling of observations closest to the treatment assignment boundary, little is known about their properties and relative merits for the analysis of BD designs. Cattaneo-Titiunik-Yu_2025_BDD-Pooling recently investigated the main empirical approaches employing pooling-based methods, offering formal identification, estimation, and inference results. Their methodological results rely on techniques from geometric measure theory (federer2014geometric), which are used to characterize how the different treatment effect estimates emerging from specifications (1) through (6) behave in large samples, that is, when $h\to 0$ as $n\to 0$. These asymptotic approximations form the basis for the continuity-based approach in standard univariate RD designs (Hahn-Todd-vanderKlaauw_2001_ECMA), and are nowadays the most common way of characterizing the properties of causal treatment effect estimators in all RD settings. This section summarizes some of the findings in Cattaneo-Titiunik-Yu_2025_BDD-Pooling.
The main challenge when studying pooling-based methods for BD designs is that the localization is done using a sequence of shrinking tubular neighborhoods covering the one-dimensional submanifold $\mathcal{B}$, as illustrated in Figure 2. An important implication is that the geometry of the treatment assignment boundary and the specific distance function used can substantially affect statistical properties. In this section, we omit most technicalities, and focus instead on the main ideas and conclusions to aid empirical researchers interpret and implement pooling-based methods.
Because the treatment assignment boundary $\mathcal{B}$ could be highly irregular, particularly in Geographic RD designs, it is necessary to restrict its geometry to establish formal properties. A minimal assumption commonly used in mathematics is to require $\mathcal{B}$ to be a rectifiable curve, that is, to assume that it has finite length. More formally, $\mathcal{B}$ is a rectifiable curve if and only if it can be reparameterized by a Lipschitz continuous function. This restriction allows for formally defining and computing integrals of functions along the one-dimensional domain $\mathcal{B}$. More importantly, under additional technical conditions, regularity on the geometry of $\mathcal{B}$ allows for those integrals to be computed over shrinking tubular neighborhoods and for limits to be well-defined, which is at the core of the results in Cattaneo-Titiunik-Yu_2025_BDD-Pooling. Under regularity conditions, the authors establish the following result:
$\lim_{h\downarrow 0}\int_{\mathcal{T}(h)}\frac{1}{h}g\Big(\frac{\mathcal{d}(\mathbf{x},\mathcal{B})}{h}\Big)m(\mathbf{x})\text{d}\mathbf{x}=\mathtt{c}_{\mathcal{B}}\cdot\int_{0}^{1}g(s)ds\cdot\int_{\mathcal{B}}\frac{m(\mathbf{x})}{\operatorname{J}_{2}\hskip-1.8063pt\mathcal{d}(\mathbf{x},\mathcal{B})}\text{d}\mathfrak{H}(\mathbf{x}),$ (9)
where $\mathcal{T}(h)=\big\{\mathbf{x}\in\mathcal{X}:\mathcal{d}(\mathbf{x},\mathcal{B})\leq h\big\}$, for each $h\geq 0$, is a tubular covering $\mathcal{B}$,
$\mathcal{d}(\mathbf{x},\mathcal{B})=\inf_{\mathbf{b}\in\mathcal{B}}\mathcal{d}(\mathbf{x},\mathbf{b})$
and $\operatorname{J}_{2}\hskip-1.8063pt\mathcal{d}(\mathbf{x},\mathcal{B})$ denotes its Jacobian, and $\mathfrak{H}(\mathbf{x})$ denotes the $1$-dimensional Hausdorff measure. The constant $\mathtt{c}_{\mathcal{B}}$ can also be characterized, but it is not important for the upcoming discussion (it can be set to $\mathtt{c}_{\mathcal{B}}=2$ for the analysis of pooling-based methods). Note that $D_{i}=\big(\mathds{1}(\mathbf{X}_{i}\in\mathcal{A}_{1})-\mathds{1}(\mathbf{X}_{i}\in\mathcal{A}_{0})\big)\cdot\mathcal{d}(\mathbf{X}_{i},\mathcal{B})$.
The integral $\int_{\mathcal{B}}\frac{m(\mathbf{x})}{\operatorname{J}_{2}\hskip-1.8063pt\mathcal{d}(\mathbf{x},\mathcal{B})}\text{d}\mathfrak{H}(\mathbf{x})$ is a natural generalization of the standard line integral, where $\text{d}\mathfrak{H}$ serves as the rigorous “length element”, enabling a robust framework for integration over complex one-dimensional domains in the plane. In particular, if $\mathcal{B}$ is piecewise linear (as in Figures 1(b) and 2), the integral can be reduced to the sum of integrals over the linear segments using a natural smooth curve parametrization and standard Riemann integration (when $m(\mathbf{x})$ is Riemann integrable). For notational simplicity, we write
$\int_{\mathcal{B}}m(\mathbf{b})\text{d}\mathbf{x}=\int_{\mathcal{B}}\frac{m(\mathbf{x})}{\operatorname{J}_{2}\hskip-1.8063pt\mathcal{d}(\mathbf{x},\mathcal{B})}\text{d}\mathfrak{H}(\mathbf{x}),$
but with the understanding that the left-hand-side integral is just notation for the rigorously defined right-hand-side integral.
The limiting integral in (9) can be used to give precise meaning to the probability limit of the treatment effect estimates from any of the pooling specifications discussed previously. However, we need to introduce standard potential outcomes notation (see, e.g., Hernan-Robins_2020_Book, for a review) to give a causal interpretation to the probability limit of those parameters. For $i=1,2,\dots,n$, let $Y_{i}(0)$ and $Y_{i}(1)$ denote the potential outcomes for unit $i$ under control and treatment assignment, respectively, and hence the observed outcome is
$Y_{i}=(1-T_{i})\cdot Y_{i}(0)+T_{i}\cdot Y_{i}(1).$
To streamline the presentation, we consider only the simplest specification (1), leading to the local difference-in-means estimator $\widehat{\tau}_{\mathtt{DIM}}$, to illustrate the methodological findings in Cattaneo-Titiunik-Yu_2025_BDD-Pooling. The other more flexible local regression specifications will enjoy similar properties.
The probability limit of $\widehat{\tau}_{\mathtt{DIM}}$ can be naturally characterized as a ratio of integrals along the boundary. As $n\to\infty$ and leveraging the logic of the law of large numbers, $\frac{1}{nh}N_{1}(h)=\frac{1}{nh}\sum_{i=1}^{n}\mathds{1}(0\leq D_{i}\leq h)/(nh)$ will be close in probability to $\frac{1}{h}\mathbb{E}[\mathds{1}(0\leq D_{i}\leq h)]$, and
$\frac{1}{h}\mathbb{E}[\mathds{1}(0\leq D_{i}\leq h)]=\int_{\mathcal{T}_{1}(h)}\frac{1}{h}g\Big(\frac{\mathcal{d}(\mathbf{x},\mathcal{B})}{h}\Big)f(\mathbf{x})\text{d}\mathbf{x}\to\int_{0}^{1}g(s)ds\cdot\int_{\mathcal{B}}f(\mathbf{x})\text{d}\mathbf{x},$
where the limit is taken as $h\to 0$, $f(\mathbf{x})$ denotes the Lebesgue density of $\mathbf{X}_{i}$, $\mathcal{T}_{1}(h)=\big\{\mathbf{x}\in\mathcal{X}:0\leq D_{i}\leq h\big\}$, and the conclusion follows by (9) upon setting $g(u)=\mathds{1}(0\leq u<1)$ and $m(\mathbf{x})=\mathds{1}(\mathbf{x}\in\mathcal{A}_{1})f(\mathbf{x})$. While $g(u)$ is redundant in this case, and $\int_{0}^{1}g(s)ds=1$, we nonetheless make it explicit to highlight that a term will appear when a non-uniform kernel is used. Similarly, $N_{0}(h)/(nh)\to_{\mathbb{P}}\int_{\mathcal{B}}f(\mathbf{x})\text{d}\mathbf{x}$. Furthermore, by the same logic, the two (properly rescaled) numerators of $\widehat{\tau}_{\mathtt{DIM}}$ will be close in probability to their expectations, where
$\frac{1}{h}\mathbb{E}[\mathds{1}(0\leq D_{i}\leq h)\cdot Y_{i}]$ $=\frac{1}{h}\mathbb{E}\Big[\mathds{1}(0\leq D_{i}\leq h)\cdot\mathbb{E}[Y_{i}(1)|D_{i}]\Big]$
$=\int_{\mathcal{T}_{1}(h)}\frac{1}{h}g\Big(\frac{\mathcal{d}(\mathbf{x},\mathcal{B})}{h}\Big)\mathbb{E}[Y_{i}(1)|\mathcal{d}(\mathbf{X}_{i},\mathcal{B})=\mathcal{d}(\mathbf{x},\mathcal{B})]f(\mathbf{x})\text{d}\mathbf{x}$
$\to\int_{0}^{1}g(s)ds\cdot\int_{\mathcal{B}}\mathbb{E}[Y_{i}(1)|\mathbf{X}_{i}=\mathbf{x}]f(\mathbf{x})\text{d}\mathbf{x},$
and, analogously,
$\frac{1}{h}\mathbb{E}[\mathds{1}(-h\leq D_{i}\leq 0)Y_{i}]\to\int_{0}^{1}g(s)ds\cdot\int_{\mathcal{B}}\mathbb{E}[Y_{i}(0)|\mathbf{X}_{i}=\mathbf{x}]f(\mathbf{x})\text{d}\mathbf{x}.$
The above calculations require precise regularity conditions and technical work, including restrictions on the geometry of $\mathcal{B}$ and the distance function $\mathcal{d}(\cdot)$ to guarantee that the expressions are well-defined and the limits are valid.
Putting the above calculations together,
$\widehat{\tau}_{\mathtt{DIM}}\to_{\mathbb{P}}\tau=\frac{\int_{\mathcal{B}}\tau(\mathbf{x})f(\mathbf{x})\text{d}\mathbf{x}}{\int_{\mathcal{B}}f(\mathbf{x})\text{d}\mathbf{x}},\qquad\tau(\mathbf{x})=\mathbb{E}[Y_{i}(1)-Y_{i}(0)|\mathbf{X}_{i}=\mathbf{x}].$
The parameter $\tau$ is called the Boundary Average Treatment Effect (BATE) by Cattaneo-Titiunik-Yu_2025_BDD-Pooling. This parameter was heuristically introduced in the education literature by Wong-Steiner-Cook_2013_JEBS, who expressed it as
$\tau=\int_{\mathcal{B}}\tau(\mathbf{x})f(\mathbf{x}|\mathbf{X}_{i}\in\mathcal{B})\text{d}\mathbf{x}=\mathbb{E}[Y_{i}(1)-Y_{i}(0)|\mathbf{X}_{i}=\mathbf{x},\mathbf{X}_{i}\in\mathcal{B}],$
using the notation $f(\mathbf{x}|\mathbf{X}_{i}\in\mathcal{B})=\frac{f(\mathbf{x})}{\int_{\mathcal{B}}f(\mathbf{x})\text{d}\mathbf{x}}$. Around the same time, Keele-Titiunik_2015_PA also discussed the BATE parameter in the context of Geographic RD designs.
The parameter $\tau$ captures a density-weighted average of average treatment effects at each point along the boundary, and thus aggregates the potentially heterogeneous treatment effects captured by $(\tau(\mathbf{x}):\mathbf{x}\in\mathcal{B})$. Cattaneo-Titiunik-Yu_2025_BDD-Location call $\tau(\mathbf{x})$ the Boundary Average Treatment Effect Curve. As discussed in Section 4, this functional causal parameter can be seen as a building block for several causal parameters of interest in BD designs.
Going beyond identification, Cattaneo-Titiunik-Yu_2025_BDD-Pooling also study estimation accuracy and uncertainty quantification for pooling-based methods. For point estimation, the authors shows that
$|\widehat{\tau}_{\mathtt{DIM}}-\tau|^{2}\approx_{\mathbb{P}}\mathsf{B}_{n}^{2}+\frac{\mathsf{V}}{nh},$
when $h\to 0$ as $n\to\infty$, and where $\approx_{\mathbb{P}}$ denotes up to higher order terms in probability, $\mathsf{B}_{n}$ denotes a leading bias term, and $\mathsf{V}$ denotes the asymptotic variance that is bounded and bounded away from zero. Characterizing the precise order of the bias is difficult without additional restrictions on the treatment assignment boundary $\mathcal{B}$: the authors show that $\mathsf{B}_{n}=O(h)$ in general. Furthermore, as in the case of standard univariate RD designs, and under additional regularity conditions, the order of the bias may be improved when $\mathbf{r}_{p}(D_{i})$ and $T_{i}\cdot\mathbf{r}_{p}(D_{i})$ are included in the local regression specification, as in (6), leading to $\mathsf{B}_{n}=O(h^{p+1})$. Another interesting feature is that the rate of decay of the variance component is the same as in one-dimensional nonparametric estimation: this improvement in convergence rate comes from the aggregation along the boundary of the bivariate nonparametric function $\tau(\mathbf{x})$. See Cattaneo-Titiunik-Yu_2025_BDD-Location and chen2025semiparametric for more discussion.
For bandwidth selection, under regularity conditions on $\mathcal{B}$ and $\mathcal{d}(\cdot)$, the bias $\mathsf{B}_{n}$ may admit a valid expansion as in the standard RD design literature. Therefore, optimal bandwidth selection procedures based on mean square error (MSE) or coverage error minimization may be used for implementation of pooling methods. See Calonico-Cattaneo-Farrell_2020_ECTJ for a review.
For distribution theory and inference, consider the usual test statistic
$\widehat{\operatorname{T}}=\frac{\widehat{\tau}_{\mathtt{DIM}}-\tau}{\sqrt{\widehat{\mathbb{V}}[\widehat{\tau}_{\mathtt{DIM}}]}},$
where $\widehat{\mathbb{V}}[\widehat{\tau}_{\mathtt{DIM}}]$ denotes any of the usual variance estimators from (local, weighted) least squares regression methods. Under regularity conditions, and if $nh\to\infty$ and $nh\mathsf{B}_{n}^{2}\to 0$, $\mathbb{P}[\widehat{\operatorname{T}}\leq u]\to\Phi(u)$, where $\Phi(u)$ denotes the standard Gaussian cumulative distribution function. Importantly, as discussed by Calonico-Cattaneo-Titiunik_2014_ECMA in the context of standard univariate RD designs, the usual confidence intervals
$\widehat{\operatorname{I}}(\alpha)=\bigg[\;\widehat{\tau}_{\mathtt{DIM}}-\mathfrak{c}_{\alpha}\cdot\sqrt{\widehat{\mathbb{V}}[\widehat{\tau}_{\mathtt{DIM}}]}\;\;,\;\;\widehat{\tau}_{\mathtt{DIM}}+\mathfrak{c}_{\alpha}\cdot\sqrt{\widehat{\mathbb{V}}[\widehat{\tau}_{\mathtt{DIM}}]}\;\bigg]$
with $\mathfrak{c}_{\alpha}=\Phi^{-1}(1-\alpha/2)$, will not be valid whenever there is local misspecification. More precisely, the validity of the usual confidence intervals $\widehat{\operatorname{I}}(\alpha)$ require the small bias condition $nh\mathsf{B}_{n}^{2}\to 0$, which is not satisfied when the bandwidth is chosen to be MSE-optimal (or is otherwise “large”). Therefore, valid uncertainty quantification depends on the choice of bandwidth and regression specification, because the bias $\mathsf{B}_{n}$ would be of order $h^{p}$, under regularity conditions on $\mathcal{B}$, $\mathcal{d}(\cdot)$, and the underlying conditional expectation functions.
If the MSE-optimal bandwidth is used, which in the context of pooling-based methods would depend on the order of the bias $\mathsf{B}_{n}$, then the small bias condition will not be satisfied, and thus the resulting inference procedures will over-reject the null hypothesis, sometimes substantially. Following Calonico-Cattaneo-Titiunik_2014_ECMA, a simple and practical solution is to employ robust bias correction. The two-step procedure is simple and intuitive: first, the pooled treatment effect is estimated using a choice of polynomial approximation $p$ and its associated MSE-optimal bandwidth; second, confidence intervals and hypothesis tests are constructed by estimating the pooled treatment effect and variance estimator using the same bandwidth choice but with a polynomial of larger order $q$. In practice, the most common choice is $p=1$ for MSE-optimal treatment effect estimation, and $q=2$ for robust bias-corrected uncertainty quantification. This inference approach has several theoretical advantages (Calonico-Cattaneo-Farrell_2018_JASA; Calonico-Cattaneo-Farrell_2022_Bernoulli), and has been validated empirically (Hyytinen-Tukiainen-etal2018_QE; DeMagalhaes-etal_2025_PA). Standard RD software for estimation and inference in univariate score settings (rdrobust, rdhte) available at https://rdpackages.github.io/ can be used directly, provided the regularity conditions are satisfied.
4 Heterogeneity and Aggregation Along the Boundary
Our review of the literature indicates that most empirical researchers using BD designs employ pooling-based methods, thereby focusing on (density-weighted) average treatment effects along the entire (or a few segments of the) boundary $\mathcal{B}$. In most implementations, the estimation begins by calculating the distance to the nearest boundary point for every observation, keeping only observations for which this distance is less than a bandwidth, and then continues by pooling all observations in a single local regression analysis, mimicking standard univariate RD methods.
Although the focus on a single scalar parameter is a convenient way of summarizing the information provided by the BD design, this approach fails to exploit its full richness. In this section, we discuss a more general approach for the analysis of BD designs that starts by studying the average treatment effect at each point on the boundary, and then uses this functional causal parameter as the building block for aggregation along the boundary, hence recovering other causal treatment effects. This section summarizes the recent results on identification, estimation and inference discussed in Cattaneo-Titiunik-Yu_2025_BDD-Distance; Cattaneo-Titiunik-Yu_2025_BDD-Location.
The starting point is to consider the Boundary Average Treatment Effect Curve (BATEC):
$\tau(\mathbf{x})=\mathbb{E}[Y_{i}(1)-Y_{i}(0)|\mathbf{X}_{i}=\mathbf{x}],\qquad\mathbf{x}\in\mathcal{B},$
which was already implicitly introduced as part of the probability limit emerging from the pooling approaches. However, the functional causal parameter $\tau(\mathbf{x})$ is of independent interest because it captures the average treatment effects at each point on the boundary. Despite capturing potentially interesting causal evidence of heterogeneous treatment effects along the assignment boundary, only a handful of empirical papers have investigated the BATEC or variations thereof. Exceptions include Keele-etal_2017_AIE, velez2019tuning, and gonzalez2021cell.
Another feature of the BATEC is that it provides the key building block for constructing other causal parameters of interest by aggregation along the boundary. For example, Cattaneo-Titiunik-Yu_2025_BDD-Location discuss the following two parameters.
  • Weighted Boundary Average Treatment Effect (WBATE):
    $\tau_{\mathtt{WBATE}}=\frac{\int_{\mathcal{B}}\tau(\mathbf{x})w(\mathbf{x})\text{d}\mathfrak{H}(\mathbf{x})}{\int_{\mathcal{B}}w(\mathbf{x})\text{d}\mathfrak{H}(\mathbf{x})},$
    where $w(\mathbf{x})$ denotes a user-chosen weighing scheme.
  • Largest Boundary Average Treatment Effect (LBATE):
    $\tau_{\mathtt{LBATE}}=\sup_{\mathbf{x}\in\mathcal{B}}\tau(\mathbf{x}).$
The WBATE parameter represents a weighted average of the (potentially heterogeneous) treatment effects $\tau(\mathbf{x})$ at each boundary point. In particular, setting $w(\mathbf{x})=f(\mathbf{x})$ recovers the BATE parameter $\tau$. In practice, other weighting schemes may also be of interest. See Reardon-Robinson_2012_JREE and Wong-Steiner-Cook_2013_JEBS for early methodological discussions, and Rischard-Branson-Miratrix-Bornn_2021_JASA for other examples considered in the context of Bayesian methods for BD designs. The insightful discussion of Qiu-Stoye_2026_BookCh--Discussion provides another application of the WBATE parameter. The LBATE parameter captures the “best” causal treatment effect along the assignment boundary, and thus can be useful for evaluating or developing targeted policies. See, for example, Andrews-Kitagawa-McCloskey_2024_QJE.
Figure 7 gives a graphical representation of the BATEC, WBATE, and LBATE in the context of the same stylized BD design used in Section 3. Because $\tau(\mathbf{x})$ is a function, estimation and inference proceeds by discretizing the boundary to construct local regression estimates for each point $\mathbf{x}\in\mathcal{B}$ on the grid of evaluation points chosen. Figure 7(a) demonstrates the idea with $40$ evenly-spaced grid points, denoted by $\mathbf{b}_{1},\ldots,\mathbf{b}_{40}$. As discussed in the upcoming subsections, distance-based and location-based methods employ these points to conduct estimation and inference, both pointwise and uniformly along the assignment boundary.
Figure 7(b) illustrates graphically the three target causal parameters $(\tau(\mathbf{x}):\mathbf{x}\in\mathcal{B})$, $\tau_{\mathtt{WBATE}}$, and $\tau_{\mathtt{LBATE}}$. The functional causal parameter $\tau(\mathbf{x})$ indicates the presence of heterogeneity along a region of the boundary, while $\tau_{\mathtt{WBATE}}$ and $\tau_{\mathtt{LBATE}}$ give two distinct notions of aggregation of those heterogeneous treatment effects. This stylized example is motivated by the SPP study by LondonoVelezRodriguezSanchez_2020_AEJ introduced above, where $\mathbf{X}_{i}=(X_{1i},X_{2i})=(\text{{SABER11}}_{i},\text{{SISBEN}}_{i})$. Re-analyzing the SPP application, Cattaneo-Titiunik-Yu_2025_BDD-Location found little evidence of heterogeneous treatment effects along the $X_{1i}$ dimension, but declining treatment effects along the $X_{2i}$ dimension. This empirical finding indicates that the average treatment effects along the boundary are roughly similar for all students with low academic performance ($\text{{SABER11}}_{i}=0$) and regardless of their wealth level ($\text{{SISBEN}}_{i}\geq 0$), while heterogeneity in treatment effects is present for the wealthier eligible students ($\text{{SISBEN}}_{i}=0$) as their academic performance increases ($\text{{SABER11}}_{i}\geq 0$).
(a) Boundary Discretization
(b) Causal Parameters
Figure 7: Heterogeneity and Aggregation along the Boundary
The discussion so far has focused on causal parameters defined along the entire $\mathcal{B}$, but in some applications researchers have considered different segments of $\mathcal{B}$. Specification (8), and its motivation from the implementation used by LondonoVelezRodriguezSanchez_2020_AEJ and others, gives one example. Employing the BATEC parameter as a building block, it is natural to define causal parameters as integrals or suprema of $\tau(\mathbf{x})$ over a specific region of $\mathcal{B}$. All the methods developed in Cattaneo-Titiunik-Yu_2025_BDD-Distance; Cattaneo-Titiunik-Yu_2025_BDD-Location immediately apply to those cases, offering estimation and inference methods for segment-specific causal parameters based on the core BATEC parameter.
4.1 Distance-Based Methods
The distance $D_{i}=\mathcal{d}(\mathbf{X}_{i},\mathcal{B})$ used by all pooling approaches is based on the distance of each observation $\mathbf{X}_{i}$ to its nearest point $\mathbf{x}\in\mathcal{B}$ on the boundary, denoted by $\mathcal{d}(\mathbf{X}_{i},\mathbf{x})$. Researchers sometimes first construct the signed distance to a pre-specified point on the boundary
$D_{i}(\mathbf{x})=\big(\mathds{1}(\mathbf{X}_{i}\in\mathcal{A}_{1})-\mathds{1}(\mathbf{X}_{i}\in\mathcal{A}_{0})\big)\cdot\mathcal{d}(\mathbf{X}_{i},\mathbf{x}),\qquad\mathbf{x}\in\mathcal{B},$
for each unit $i=1,\ldots,n$, and then use this distance to construct a tubular covering $\mathcal{B}$ to implement pooling methods. One example of this approach is ehrlich2018persistent. While the resulting tubular will often be geometrically different from $\mathcal{T}(h)=\{\mathbf{x}\in\mathcal{X}:\mathcal{d}(\mathbf{x},\mathcal{B})\leq h\}$, all the ideas and results discussed in Section 3 continue to apply.
However, the intermediate boundary-point-specific signed distance $(D_{i}(\mathbf{x}):\mathbf{x}\in\mathcal{B})$ for each unit can also be used directly to estimate and conduct inference for the BATEC. The idea is to view the outcomes and the boundary-point-specific scalar distance variables, $(Y_{1},D_{1}(\mathbf{x})),\ldots,(Y_{n},D_{n}(\mathbf{x}))$, as a one dimensional RD design for each point $\mathbf{x}\in\mathcal{B}$. For example, this approach was used by Keele-Titiunik_2015_PA and velez2019tuning to study the effect of TV exposure on voter turnout, using reception and media market boundaries that determine whether citizens are exposed to specific television programming. In these examples, each unit’s score $\mathbf{X}_{i}$ is a latitude-longitude pair that determines their place of residence, the outcome $Y_{i}$ is whether the individual turned out to vote, and the boundary $\mathcal{B}$ is the media market or reception boundary that separates the treated from the control region. In Keele-Titiunik_2015_PA, the treatment is exposure to presidential candidate television ads during the presidential election, and in velez2019tuning the treatment is exposure to Spanish-language television. In both cases, researchers first selected a grid $\mathbf{b}_{1},\mathbf{b}_{2},\ldots\mathbf{b}_{J}$ on the boundary (e.g. Figure 7(a)), and then used the local regression specification:
$\mathtt{reg}\quad Y_{i}\quad\mathtt{on}\quad 1,\;T_{i},\;\mathbf{r}_{p}(D_{i}(\mathbf{b}_{j})),\;T_{i}\cdot\mathbf{r}_{p}(D_{i}(\mathbf{b}_{j}))\qquad[\;\mathds{1}(|D_{i}(\mathbf{b}_{j})|\leq h)\;],$ (10)
for each $j=1,\ldots,J$. The key difference is that estimation is conducted with localization for each evaluation point $\mathbf{b}_{j}$ on the boundary, as opposed to localization to the entire boundary as in specifications (1) through (8). Therefore, the estimated coefficient accompanying $T_{i}$ gives a point estimate of $\tau(\mathbf{b}_{j})$ for each $j=1,\ldots,J$
Cattaneo-Titiunik-Yu_2025_BDD-Distance study the large sample properties of the distance-based estimation procedure for BATEC obtained from specification (10). They give necessary and/or sufficient conditions for identification, estimation, and inference in large samples, both pointwise and uniformly along the boundary. Their identification result requires minimal regularity of the boundary $\mathcal{B}$, the distance function $\mathcal{d}(\cdot)$, and the underlying data generating process, building on continuity-based identification in standard univariate RD designs (Hahn-Todd-vanderKlaauw_2001_ECMA). Cattaneo-Titiunik-Yu_2025_BDD-Distance also show that the bias of the distance-based point estimator will depend on the smoothness of the assignment boundary, while the variance of the estimator will be of order $(nh^{2})^{-1}$. Putting these two results together, it follows that bandwidth selection methods for standard univariate RD designs will either minimize the mean square error of the estimator or deliver undersmoothing, depending on the geometry of the assignment boundary and other assumptions. For inference, the authors establish asymptotic validity of confidence intervals and confidence bands, provided the bias of the distance-based point estimator is small enough (which, in turn, crucially depends on the geometry of $\mathcal{B}$).
4.2 Location-Based Methods
While the distance-based estimator constructed using specification (10) provides a natural bridge between the pooling methods and heterogeneity analysis, Cattaneo-Titiunik-Yu_2025_BDD-Distance demonstrate that this approach would require a smooth assignment boundary for higher-order bias reduction. Otherwise, the distance-based approach will generate a consistent estimator with possibly a high bias, and therefore require a small bandwidth for estimation and inference validity. A solution to this problem is to use local bivariate regression methods.
Cattaneo-Titiunik-Yu_2025_BDD-Location study the properties of local regression treatment effect estimators directly employing the location information encoded in $\mathbf{X}_{i}$ for each unit. The authors develop pointwise and uniform estimation and inference methods for both the BATEC and transformations thereof, such as the WBATE and LBATE. Their local regression specification takes the form:
$\mathtt{reg}\quad Y_{i}\quad\mathtt{on}\quad 1,\;T_{i},\;\mathbf{r}_{p}(\mathbf{X}_{i}-\mathbf{b}_{j}),\;T_{i}\cdot\mathbf{r}_{p}(\mathbf{X}_{i}-\mathbf{b}_{j})\qquad[\;K\big(\tfrac{\mathbf{X}_{i}-\mathbf{b}_{j}}{h}\big)\;],$ (11)
where $\mathbf{b}_{1},\ldots,\mathbf{b}_{J}$ is the grid of points on the boundary (e.g., Figure 7(a)). This local polynomial specification crucially depends on $\mathbf{X}_{i}$ directly via the $p$th order polynomial expansion $\mathbf{r}_{p}(\mathbf{X}_{i}-\mathbf{x})$. The weights are determined by the bivariate kernel function $K(u_{1},u_{2})$, and the localization to each point $\mathbf{b}_{j}$ continues to be is controlled by the bandwidth $h$. In particular, $K\big(\frac{\mathbf{X}_{i}-\mathbf{b}_{j}}{h}\big)=\mathds{1}(\|\mathbf{X}_{i}-\mathbf{b}_{j}\|\leq h)=\mathds{1}(|D_{i}(\mathbf{b}_{j})|\leq h)$ is a valid choice, demonstrating an interesting connection with distance-based methods as implemented via specification (10). The estimated coefficient accompanying $T_{i}$, denoted by $\widehat{\tau}(\mathbf{b}_{j})$, gives the location-based point estimate of $\tau(\mathbf{b}_{j})$, for each $j=1,\ldots,J$.
In the location-based setting, identification of $\tau(\mathbf{x})$ for $\mathbf{x}\in\mathcal{B}$ follows directly from a generalization of continuity-based identification results in the standard univariate RD design (see Hahn-Todd-vanderKlaauw_2001_ECMA; Papay-Willett-Murnane_2011_JoE; Reardon-Robinson_2012_JREE; Wong-Steiner-Cook_2013_JEBS; Keele-Titiunik_2015_PA; Cattaneo-Keele-Titiunik-VazquezBare_2016_JOP). Under minimal regularity conditions on the boundary and data generating process, Cattaneo-Titiunik-Yu_2025_BDD-Location establishes (pointwise and) uniform convergence rates of the form $\sup_{\mathbf{x}\in\mathcal{B}}|\widehat{\tau}(\mathbf{x})-\tau(\mathbf{x})|^{2}=O_{\mathbb{P}}(\tfrac{\log(n)}{nh^{2}}+h^{2(p+1)})$, demonstrating that the location-based estimator does not require stringent smoothness restrictions on the boundary $\mathcal{B}$ to achieve higher order debiasing. Furthermore, the authors establish pointwise and integrated mean square error expansions for the estimator, and develop MSE-optimal bandwidth selection methods. By combining these results, they develop optimal point estimation of the BATEC, both pointwise and uniform, based on specification (11).
For uncertainty quantification, Cattaneo-Titiunik-Yu_2025_BDD-Location consider confidence intervals and confidence bands of the form:
$\widehat{\operatorname{I}}(\mathbf{x},\alpha)=\bigg[\;\widehat{\tau}(\mathbf{x})-\mathfrak{c}_{\alpha}\cdot\sqrt{\widehat{\mathbb{V}}[\widehat{\tau}(\mathbf{x})]}\;\;,\;\;\widehat{\tau}(\mathbf{x})+\mathfrak{c}_{\alpha}\cdot\sqrt{\widehat{\mathbb{V}}[\widehat{\tau}(\mathbf{x})]}\;\bigg],\qquad\mathbf{x}\in\mathcal{B},$
where $\widehat{\mathbb{V}}[\widehat{\tau}(\mathbf{x})]$ denotes any of the usual variance estimators from (weighted) least squares regression methods, and $\mathfrak{c}_{\alpha}$ is the quantile used depending on inferential goal. For confidence intervals, the usual Gaussian quantile is asymptotically valid (i.e., $\mathfrak{c}_{\alpha}=\Phi^{-1}(1-\alpha/2)$), while for confidence bands, a new (larger) quantile needs to be constructed to capture the joint uncertainty features of the estimators $\widehat{\tau}(\mathbf{x})$ for different values along the boundary $\mathcal{B}$. As in the case of pooling-based and distance-based methods, a key condition needed for validity of these confidence intervals and bands is the small (misspecification) bias property. Cattaneo-Titiunik-Yu_2025_BDD-Location address this issue by relying on robust bias correction (Calonico-Cattaneo-Titiunik_2014_ECMA; Calonico-Cattaneo-Farrell_2022_Bernoulli), a construction that proceeds in two steps.
  • Step 1. For a choice $p$ (usually $p=1$), the MSE-optimal bandwidth is computed, and then for that choice of bandwidth the point estimator $\widehat{\tau}(\mathbf{x})$ is computed using specification (11). As a result, $\widehat{\tau}(\mathbf{x})$ is an asymptotically MSE-optimal point estimator of $\tau(\mathbf{x})$.
  • Step 2. For a choice $q>p$ (usually $q=2$), the point estimator $\widehat{\tau}(\mathbf{x})$ and variance estimator $\widehat{\mathbb{V}}[\widehat{\tau}(\mathbf{x})]$ are computed using specification (11) with the same bandwidth used in Step 1. As a result, inference procedures are robust bias-corrected, and hence asymptotically valid.
Cattaneo-Titiunik-Yu_2025_BDD-Location give technical and computational details, which we omit to conserve space, and because they do not fundamentally change the core methodological ideas.
Finally, $\widehat{\tau}(\mathbf{x})$ can also be used to estimate and conduct inference for WBATE and LBATE via plug-in methods. Cattaneo-Titiunik-Yu_2025_BDD-Location provide sufficient conditions and study the validity of such methods. In particular, the WBATE corresponds to an integral over a submanifold of the nonparametric estimator $\widehat{\tau}(\mathbf{x})$. See chen2025semiparametric for related theoretical results when employing series approximations instead of local polynomial regression as commonly done in RD settings.
The general-purpose R package rd2d implement both distance-based and location-based methods (https://rdpackages.github.io/). See Cattaneo-Titiunik-Yu_2025_rd2d for more discussion. Figure 8 gives a stylized example of the type of results that can be obtained via distance-based and location-based methods, building on the same setup as in Figure 1. These figures are motivated by empirical work using the SPP application (LondonoVelezRodriguezSanchez_2020_AEJ).
Figure 8: Estimation and Inference for BATEC
5 Recommendations for Practice
Our review of the empirical literature employing BD designs showed that the overwhelming majority of studies focus exclusively on the identification, estimation and inference for the BATE using pooling-based methods (Section 3). This scalar parameter captures a density-weighted average of potentially heterogeneous treatment effects along the assignment boundary. While modern empirical work has recognized the importance of localization and flexible local parametrization, the current literature often exhibits three key limitations: (i) local regression specifications frequently omit the interaction between the treatment assignment variable $T_{i}$ and the flexible regression terms $\mathbf{r}_{p}(D_{i})$ or $\mathbf{r}_{p}(\mathbf{X}_{i})$, risking significant misspecification bias; (ii) the bandwidth $h$ is often chosen in an ad hoc manner; and (iii) the role of misspecification bias from a “large” bandwidth, such as an MSE-optimal choice, is often ignored, which can lead to over-rejection of the null hypothesis. Cattaneo-Titiunik-Yu_2025_BDD-Pooling provide a formal study of pooling-based methods that explicitly addresses these limitations. Moving forward, we recommend that empirical researchers implementing pooling-based methods always use specifications (6) or (7), select an MSE-optimal bandwidth, and conducting inference using robust bias-corrected methods.
More broadly, as noted by Cattaneo-Titiunik-Yu_2025_BDD-Distance; Cattaneo-Titiunik-Yu_2025_BDD-Location, a more general approach to analyzing and interpreting BD designs is to first focus on the BATEC parameter, and then consider other aggregate causal treatment effects as transformations thereof. As summarized in Section 4, distance-based and location-based methods capture the full richness of the BD design by estimating average treatment effects at each point along the assignment boundary. These methods also recover the BATE parameter as a special case, in addition to other interesting causal parameters such as the LBATE. These methods can be implemented using the general-purpose R package rd2d available at https://rdpackages.github.io/, and discussed in Cattaneo-Titiunik-Yu_2025_rd2d. In particular, location-based methods are the most robust and general approach; we therefore recommend them for future empirical work leveraging BD designs.
Selected Empirical Articles Using The Boundary Discontinuity Design
Article $Y_{i}$ $T_{i}$ $\mathbf{X}_{i}$ Distance Specifications Bandwidth Heter
[…]
Article $Y_{i}$ $T_{i}$ $\mathbf{X}_{i}$ Distance Specifications Bandwidth Heter
[…]
holmes1998effect Manufacturing activity Right-to-work policies Lat, Lon Euclidean (1), (2), (6) Manual Yes
black1999better House prices School quality Lat, Lon Euclidean (1), (2) Manual No
kane2003school House prices School quality Lat, Lon Euclidean (1), (2) Manual No
kane2006school Housing prices School quality Lat, Lon Euclidean (1), (2) Manual No
bayer2007unified House prices School quality Lat, Lon Euclidean (1), (2) Manual No
lalive2008extended Unemployment duration Unemployment benefits Lat, Lon Driving time (6) Manual No
dell2010persistent Consumption, childhood stunting Forced mining labor Lat, Lon Euclidean (3) MSE No
ou2010leave High school dropout Failure of high school exit exam Mathematics score, Language score Euclidean (8) Manual No
dube2010minimum Employment Increase in minimum wage Lat, Lon County adjacency (1), (2) Manual No
grout2011land Property values Land-use regulations Lat, Lon Euclidean (8) Manual Yes
robinson2011evaluating English proficiency English proficiency reclassification Scores on five language tests Euclidean Binding score MSE No
eugster2011demand Demand for social insurance Culture Lat, Lon Driving kilometers (6) Manual No
hinnerich2014democracy Political regimes Income redistribution Population in two separate years Euclidean (8) MSE No
ferwerda2014political Resistance activity Devolution of governing authority Lat, Lon Euclidean (6) MSE, Manual No
snider2015barriers Airline industry outcomes Access to airport facilities granted to new airlines Share of passengers in top two carriers in origin and destination city (1), (11) MSE, Manual Yes
barone2015telecracy Vote shares Switch from analog to digital TV Lat, Lon Euclidean (5) Manual No
michalopoulos2016long Civil conflict Ethnic partitioning Lat, Lon Euclidean (1), (2) Manual No
egger2015impact Foreign investment Controlled foreign company rules Tax rate, Income measures NA (7) $L=1$ MSE No
Keele-Titiunik_2015_PA Voter turnout Political TV advertisement Lat, Lon Chordal (10) MSE Yes
macdonald2016effect Crime Policing Lat, Lon Euclidean (6) MSE No
evans2017smart STEM major Financial incentive GPA, Family contribution score Euclidean (8) MSE No
eugster2017culture Unemployment Culture Lat, Lon Road distance (6) Manual No
kumar2018restrictions Mortgage default Limits on home equity borrowing Lat, Lon Euclidean (4), (3) Manual No
[…]
dell2018nation Public goods, political attitudes War strategies Lat, Lon Euclidean (3) Manual No
clinton2018politics Voter registration and turnout Medicaid expansion Lat, Lon Euclidean (6) Manual No
spenkuch2018political Voter turnout Television advertisement Lat, Lon Euclidean (2), (6) MSE, Manual No
de2018agents Party vote shares Political alignment Lat, Lon Euclidean (3) Manual No
eugster2019culture Tax competition Culture Lat, Lon Road (1) in DID Manual No
giuntella2019sunset Sleep duration, Health outcomes Timing of natural light Lat, Lon Euclidean (6) MSE, Manual No
johnson2019effects High School graduation English learner classification Scores on two language tests Euclidean Binding score MSE No
velez2019tuning Voter turnout Spanish-language television station Lat, Lon Euclidean (1), (2), (10) MSE, Manual Yes
dupraz2019french Schooling Colonial legacies Lat, Lon Euclidean (6), (3) MSE No
moscona2020segmentary Conflict Segmentary lineage organization Lat, Lon Euclidean (6) Manual No
ambrus2020loss Property value Cholera epidemic Lat, Lon Walking (6) $L=1$ MSE, Manual No
dell2020development Schools, education levels Sugar cultivation for Dutch cultivation system Lat, Lon Euclidean (3) Manual No
he2020watering Pollution Water quality monitoring Lat, Lon Euclidean (6) $L=1$ MSE No or NA
ito2020willingness Willingness to pay for clean air Lat, Lon Euclidean, road (6) MSE, Manual No
wuepper2020countriesErosion Soil erosion Country borders Lat, Lon Euclidean (6) MSE No
albertus2020land Conflict Land reform Lat, Lon Euclidean (6) MSE, Manual No
wuepper2020countriesPollution Crop yield gaps, nitrogen pollution Country borders Lat, Lon Euclidean (6) MSE No
schafer2020time Voter turnout Time zones Lat, Lon Chordal (6) MSE, Manual No
letsa2020mechanisms Attitudes towards local power Colonial legacies Lat, Lon Euclidean (3) Manual No
LondonoVelezRodriguezSanchez_2020_AEJ College enrollment Financial subsidy High school exit exam score, wealth index Euclidean (8) MSE No
gonzalez2021cell Election fraud Cell phone coverage Lat, Lon Euclidean (6), (7) MSE Yes
ayres2021environmental Land value Adjudication of groundwater rights Lat, Lon Euclidean (6) $L=1$ MSE No
laliberte2021long School quality Educational achievement Lat, Lon Euclidean (6) MSE No
aaronson2021effects Characteristics of urban neighborhoods Red lining maps Lat, Lon Euclidean (2) in DID Manual No
appau2021long Agricultural productivity Bombing intensity Lat, Lon Absolute difference between centroid and 17th parallel north latitude (4) (first stage) Manual No
lowes2021concessions Education, health, wealth Concession to extract rubber Lat, Lon Euclidean (6) MSE, Manual No
sides2022effect Election outcomes Television advertisement Lat, Lon Euclidean (2) Manual No
zheng2022valuation House prices Charter school entry Lat, Lon Euclidean Event study in sample near boundary Manual No
dehdari2022origins Regional identity Actions of nation states Lat, Lon Euclidean (6) MSE, Manual No
mendez2022multinationals Living standards Land concession to multinational firm Lat, Lon Euclidean (3) Manual No
jones2022factor Profits, irrigation efficiency Irrigation Lat, Lon Euclidean (6) Manual No
henn2023complements State capacity Remoteness from administrative headquarters Lat, Lon Euclidean (6) Manual No
jones2022effects Enrollment, graduation Tuition scholarship High school GPA, SAT/ACT score NA (7) Manual No
dai2022effects Lifestyle outcomes Hypertension diagnosis Systolic and Diastolic blood pressure Euclidean (8) MSE No
salti2022impact Economic outcomes Cash assistance Two scores in means testing formula NA (8) MSE No
castro2022effect Teacher retention, student learning Recruitment bonus Population, distance to capital Euclidean (8) MSE No
moussa2022impact Children health outcomes Cash assistance Two scores in means testing formula NA (8) MSE No
mangonnet2022playing Protected area designation Political alignment Lat, Lon Euclidean (2) Manual No
baragwanath2023collective Forest growth Property rights Lat, Lon Euclidean (6) MSE No
larsen2024long Educational attainment Grade retention Mathematics and English Language Arts test scores Euclidean (8) MSE No
woller2023cost Vote buying Cost of voting Lat, Lon Euclidean (4) Manual No
paulsen2023foundations Voter turnout, economic outcomes State funding for common schools Lat, Lon Euclidean (3) Manual No
prillaman2023strength Political participation Participation in women-only credit groups Lat, Lon Euclidean (2), (3) Manual No
mcalexander2023international Rebel actions UN partition line of Palestine Lat, Lon Euclidean (6) $L=1$ MSE No
murphy2023dual Special education placement English learner designation Various scores of language assessment NA Binding score MSE No
bjerre2024attendance School enrollment School change Lat, Lon Euclidean (2) in DID Manual No
kampfen2024heterogeneous Blood pressure Blood pressure diagnosis Systolic and Diastolic blood pressure NA Binding score MSE No
wuepper2024public Deforestation Country borders Lat, Lon Euclidean (6) MSE No
cox2024bound Voter mobilization Social networks Lat, Lon Fastest driving (1) NA No
[…]
grasse2024state Poverty, economic development Mass repression Lat, Lon Euclidean (5) Manual No
jardim2024local Wages, employment Increase in minimum wage Lat, Lon Driving time (1), (2) Manual No
ring2025wealth Savings Wealth tax assessment Lat, Lon Euclidean (6) Manual No
yamagishi2025persistent Land price Buraku neighborhoods Lat, Lon Euclidean (1), (6) $L=1$ MSE, Manual No
loumeau2025regional Commuting flows, residential decisions Departmental borders Lat, Lon Euclidean (5) MSE No
boix2025political Jewish National identity Political emancipation Lat, Lon Euclidean (6) MSE No
muller2025building Share of predominant ethnic group Administrative borders Lat, Lon Euclidean (6) Manual No